# Hybrid quantum cycle generative adversarial network for small molecule generation

### v1

@@ II.1 HQ-MolGAN
The investigation into the efficacy of the Hybrid Generator within the MolGAN framework commences with a comparative analysis focusing on the chemical properties of the generated molecular samples. Specifically, this analysis evaluates the logP and the QED scores. By contrasting the logP and QED scores yielded by the molecules generated from each model, we aim to quantify and elucidate the impact of the Hybrid Generator’s integration on the model’s performance in generating chemically viable and optimally structured molecules.
@@
A potential explanation for the discontinuous score trends observed in the HQ-MolGAN could be attributed to its limited validation set size.
Nevertheless, the HQ-MolGAN capable to generate molecular samples with competitive scores in terms of Drug-likeness, Synthesizability, and Solubility, as depicted in Fig. 4(a). As shown in Table 3, the HQ-MolGAN-VVRQ model trained on the QM9 dataset achieves a LogP score of $0.84$, the HQ-MolGAN-EFQ model trained on the PC9 dataset results in a QED score of $0.62$, and the HQ-MolGAN-EFQ model trained on both datasets attains an SA score of $0.84$.
@@
During our experiments, a notable limitation of the Hybrid models was observed, characterized by a tendency of the generator to gravitate towards a “high entropy state”. This phenomenon is illustrated in Fig. 4(b)). In this state, the generator predominantly produces molecular structures that are either bare, unbound atoms or a collection of disconnected small molecules. Intriguingly, despite their simplistic and fragmented nature, these structures are often assigned high scores in terms of LogP and SA by the RDKit library within the Reward component. This paradoxical scoring poses a challenge to the model’s reliability in generating chemically meaningful and complex molecules.
This observation indicates a critical issue in the mapping process of the generator. Essentially, a variety of noise samples $Z$ are being mapped to a limited and similar region in the chemical space $Y^{\prime}$, resulting in repetitive and high-entropy molecular samples. Such a mapping significantly diminishes the expressivity of the Generator, constraining its ability to generate a diverse range of molecular structures.
Figure 5: Comparison of combined losses during training on the QM9 dataset. (a): Cycle-MolGAN versus MolGAN models, Cycle-MolGAN has a more stable training process compared to the MolGAN. (b): HQ-Cycle-MolGAN-EFQ versus Hybrid-MolGAN-EFQ, no significant impact of Cycle component on Loss curve observed.
Figure 6: (a) Graph of combined losses of MolGAN and HQ-Cycle Component MolGAN models trained on QM9 dataset. (b): Graph of combined losses of HQ-Cycle-MolGAN-EFQ and HQ-MolGAN-VVRQ models trained on QM9 dataset.
@@
Table 2: MolGAN and HQ-MolGAN with classic Cycle component
Table 3: MolGAN and HQ-MolGAN with Hybrid-Quantum Cycle component
@@
HQ-MolGAN-VVRQ VVRQ (Both Datasets) $68.8$ 11.5 0.98 $0.52$ 0.63 0.75
@@
Cycle-MolGAN (Both Datasets) $\mathbf{93.9}$ $6.52$ $0.99$ $\mathbf{0.49}$ $0.25$ $0.78$
HQ-Cycle-MolGAN-VVRQ (QM9) $60.4$ 8.7 0.94 $0.53$ $0.38$ 0.61
HQ-Cycle-MolGAN-VVRQ (PC9) $67.8$ 9.1 0.94 $0.53$ $\mathbf{0.50}$ 0.93
@@
Table 2: MolGAN and HQ-MolGAN with classic Cycle component
Table 3: MolGAN and HQ-MolGAN with Hybrid-Quantum Cycle component
@@ II.2 HQ-Cycle-MolGAN
Prior to assessing the effect of the Cycle component on the training of HQ-MolGAN, it is essential to first examine its impact on the conventional MolGAN framework. As indicated in Fig. 5(a), the incorporation of the Cycle component into MolGAN (termed Cycle-MolGAN) results in a more stable training process compared to the ordinary MolGAN model. This stability significantly enhances the quality of the generated molecular samples, as reflected in their improved Diversity scores (Table 3, 3). Furthermore, the Cycle component contributes to the generation of more complex and “bounded” molecular structures, indicating a higher degree of chemical realism.
In the context of HQ-MolGAN, while the integration of the Cycle component does not markedly alter the Loss curve as depicted in Fig. 5(b), its influence is evident in the improved key metric scores of the final HQ-Cycle-MolGAN models. Whether trained on the PC9 dataset or on a combination of datasets, the HQ-Cycle-MolGAN demonstrates superior performance in key metrics, as shown in Table 3. The stabilizing properties of the Cycle component aid the model in consistently generating “bounded” molecular samples. Notably, both the VVRQ and EFQ variants of HQ-MolGAN achieve significant scores in terms of LogP ($\mathbf{0.93}$ and $\mathbf{0.94}$) and QED.
Additionally, an analysis of Tables 3 and 3 reveals that models equipped with the Cycle component are capable of producing a greater number of unique samples. This finding aligns with the intended objective of the Cycle component, which is to navigate through the “high entropy state” and enhance the diversity and uniqueness of the molecular samples generated by the model.
@@ II.3 Hybrid-Quantum Cycle MolGAN
In the third and final stage of our experimental series, we focused on evaluating the impact of the Hybrid-Quantum Cycle Component on both the classical MolGAN and the HQ-MolGAN architectures.
Fig. 6 presents a comparative analysis of the Generator losses between MolGAN and HQ-MolGAN models with and without the Hybrid-Quantum Cycle Component. According to the data presented in Table 3, the incorporation of the Hybrid-Quantum Cycle Component does not result in significant improvements in most of the desired metrics, except for a notable LogP score of $0.95$ achieved by the HQ-Cycle-MolGAN-VVRQ model trained on both datasets.
This absence of a marked enhancement in performance metrics for models incorporating the HQ-Cycle component, as compared to their counterparts without it, could potentially be attributed to an insufficient number of training iterations. This hypothesis is supported by the observations from Fig. 6, where the training losses of MolGAN and MolGAN with the HQ-Cycle component exhibit minimal divergence, suggesting that extended training might be necessary for realizing the full potential of the HQ-Cycle Component.
@@ III Discussion
To enhance the classical MolGAN, we introduce two solutions: the incorporation of VQCs as the initial layer of the generator and the utilization of a cycle component to restore the original data from the graph representation of the generated molecule.
Our empirical results substantiate the merit of diversifying training datasets, not limiting to the QM9 dataset alone but also incorporating the PC9 dataset or a combination of both. Notably, the HQ-MolGAN model, with their Generator’s layers scaled down by half and trained for a quarter fewer iterations, has outperformed the classical MolGAN model decao2022molgan and its hybrid quantum analogues Li_Topaloglu_Ghosh_2021 ; Kao_2023 across key chemical metrics: QED, logP, SA and uniqueness.
The introduction of the HQ-Cycle-MolGAN, and especially its variant with the HQ-Cycle component, marks a significant advancement in the training of hybrid quantum models. These models not only enhance the desired uniqueness score but also effectively mitigate the occurrence of the “high entropy state”, a notable challenge in molecular generation tasks. Consequently, these models hold substantial promise for applications in the domain of small drug compound design, both for commercial and scientific purposes within the pharmacology.
@@
Looking ahead, we see a potential for hybrid quantum machine learning models to further advance the field of molecule generation using hybrid quantum models. We plan to delve deeper into refining the model architecture, particularly focusing on optimizing the balance between the quantum and classical components. This involves experimenting with different configurations and parameters to enhance the overall efficiency and accuracy of the models. Another critical avenue we intend to pursue is the expansion of our training datasets. By incorporating a broader range of chemical compounds and molecular structures, we aim to increase the diversity and representativeness of our models. This expansion is expected to improve the models’ generalization capabilities and their applicability. Through these focused research efforts, we aspire to contribute significantly to the advancement of hybrid quantum computing in drug discovery, ultimately aiding in the development of more effective and innovative therapeutic solutions.

### v2

@@ II.1 HQ-MolGAN
We start the investigation into the efficacy of the Hybrid Generator within the MolGAN framework with a comparative analysis focusing on the chemical properties of the generated molecular samples. Specifically, this analysis evaluates the logP and QED scores. By contrasting the logP and QED scores yielded by the molecules generated from each model, we aim to quantify and elucidate the impact of the Hybrid Generator’s integration on the model’s performance in generating chemically viable and optimally structured molecules.
@@
A potential explanation for the discontinuous score trends observed in the HQ-MolGAN could be attributed to its limited validation set size. Nevertheless, the HQ-MolGAN can generate molecular samples with competitive scores in terms of Drug-likeness, Synthesizability, and Solubility, as depicted in Fig. 4(a). As shown in Table 4, the HQ-MolGAN-VVRQ model trained on the QM9 dataset achieves a LogP score of $0.84$, the HQ-MolGAN-EFQ model trained on the PC9 dataset results in a QED score of $0.62$, and the HQ-MolGAN-EFQ model trained on both datasets attains an SA score of $0.84$.
@@
During our experiments, we observed a notable limitation of the hybrid models, characterized by the generator’s tendency to gravitate towards a ”high entropy state.” This phenomenon is illustrated in Fig. 4(b). In this state, the generator predominantly produces molecular structures that are either bare, unbound atoms or a collection of disconnected small molecules. Intriguingly, despite their simplistic and fragmented nature, these structures are often assigned high scores in terms of LogP and SA by the RDKit library within the Reward component. This paradoxical scoring poses a challenge to the model’s reliability in generating chemically meaningful and complex molecules.
This observation indicates a critical issue in the generator’s mapping process. Essentially, various noise samples $Z$ are mapped to a limited and similar region in the chemical space $Y^{\prime}$, resulting in repetitive and high-entropy molecular samples. Such a mapping significantly diminishes the generator’s expressivity, constraining its ability to generate a diverse range of molecular structures.
Figure 5: Figures (a)-(d) present a comparison of the combined losses during training on the QM9 dataset: (a) MolGAN and Cycle-MolGAN. Cycle-MolGAN has a more stable training process compared to the MolGAN. (b) HQ-Cycle-MolGAN-EFQ and HQ-MolGAN-VVRQ. (c) MolGAN an HQ-Cycle-Component MolGAN. (d) HQ-Cycle-MolGAN-EFQ versus Hybrid-MolGAN-EFQ. No significant impact of the Cycle component on the loss curve is observed. Figures (e)-(f) present charts for the IBM Brisbane execution: (e) Graph of the relative errors of the simulators’ probability matrices with respect to the number of shots. (f) Comparison of the probabilities generated by the noisy and ideal simulators using $2\times 10^{5}$ shots.
@@
HQ-MolGAN-VVRQ (Both Datasets) $68.8$ 11.5 0.98 $0.52$ 0.63 0.75
@@
Table 2: MolGAN and HQ-MolGAN with classic Cycle component
@@
Table 3: MolGAN and HQ-MolGAN with Hybrid-Quantum Cycle component
@@
Cycle-MolGAN (Both Datasets) $\mathbf{93.9}$ $6.52$ $0.99$ $0.49$ $0.25$ $0.78$
HQ-Cycle-MolGAN-VVRQ (QM9) $60.4$ 8.7 0.94 $\mathbf{0.53}$ $0.38$ 0.61
HQ-Cycle-MolGAN-VVRQ (PC9) $67.8$ 9.1 0.94 $\mathbf{0.53}$ $\mathbf{0.50}$ 0.93
@@
Table 4: HQ-MolGAN-VVRQ forward pass on the noisy (IBM Brisbane) and ideal (Qiskit) simulators using $2\times 10^{5}$ shots budget.
[…]
Noisy HQ-MolGAN-VVRQ (QM9) $\mathbf{80.0}$ $\mathbf{6.52}$ $\mathbf{0.97}$ $\mathbf{0.44}$ $\mathbf{0.23}$ $\mathbf{0.76}$
Ideal HQ-MolGAN-VVRQ (QM9) $\mathbf{80.0}$ $6.51$ $\mathbf{0.97}$ $\mathbf{0.44}$ $\mathbf{0.23}$ $0.75$
@@ II.2 HQ-Cycle-MolGAN
Prior to assessing the effect of the Cycle component on the training of HQ-MolGAN, it is essential to first examine its impact on the conventional MolGAN framework. As indicated in Fig. 5(a), the incorporation of the Cycle component into MolGAN (termed Cycle-MolGAN) results in a more stable training process compared to the ordinary MolGAN model. This stability significantly enhances the quality of the generated molecular samples, as reflected in their improved Uniqueness scores (Table 4, 4). Furthermore, the Cycle component contributes to the generation of more complex and “bounded” molecular structures, indicating a higher degree of chemical realism (Fig. 4(c, d)).
In the context of HQ-MolGAN, while the integration of the Cycle component does not markedly alter the loss curve as depicted in Fig. 5(b), its influence is evident in the improved key metric scores of the final HQ-Cycle-MolGAN models. Whether trained on the PC9 dataset or a combination of datasets, the HQ-Cycle-MolGAN demonstrates superior performance in key metrics, as shown in Table 4. The stabilizing properties of the Cycle component aid the model in consistently generating “bounded” molecular samples. Notably, both the VVRQ and EFQ variants of HQ-MolGAN achieve significant scores in terms of LogP ($\mathbf{0.93}$ and $\mathbf{0.94}$) and QED.
Additionally, an analysis of Tables 4 and 4 reveals that models equipped with the Cycle component are capable of producing a greater number of unique samples. This finding aligns with the intended objective of the Cycle component, which is to navigate through the “high entropy state” and enhance the diversity and uniqueness of the molecular samples generated by the model. The addition of Cycle components provides a more stable (even smoother) training process for classical MolGAN. In terms of HQ-MolGAN, it rapidly increases its isomorphic properties that significantly help to get through a “high entropy state” during training. This increase in objectivity can be seen in the increase in uniqueness scores of cycle models.
@@ II.3 Hybrid-Quantum Cycle MolGAN
In the third stage of our experimental series, we focused on evaluating the impact of the Hybrid-Quantum Cycle Component on both the classical MolGAN and the HQ-MolGAN architectures. In our simulations, the VVRQ generator uses a quantum circuit with $3\times 8\times 3$ parameters. In the case of the QDI layer in the HQ-Cycle component, the quantum circuit has $8\times 8$ parameters.
Fig. 5(c-d) presents a comparative analysis of the Generator losses between MolGAN and HQ-MolGAN models with and without the Hybrid-Quantum Cycle Component. According to the data presented in Table 4, the incorporation of the Hybrid-Quantum Cycle Component does not result in significant improvements in most of the desired metrics, except for a notable LogP score of $0.95$ achieved by the HQ-Cycle-MolGAN-VVRQ model trained on both datasets.
This absence of a marked enhancement in performance metrics for models incorporating the HQ-Cycle component, as compared to their counterparts without it, could potentially be attributed to an insufficient number of training iterations. This hypothesis is supported by the observations from Fig. 5(c), where the training losses of HQ-MolGAN and HQ-MolGAN with the HQ-Cycle component exhibit minimal divergence, suggesting that extended training might be necessary for realizing the full potential of the HQ-Cycle Component.
@@
II.4 Execution on simulators of noisy quantum devices
In the last section of our numerical experiments, we explore the potential of executing the HQ-MolGAN-VVRQ model on quantum devices. For that exploration we generated molecular samples using two IBM simulators: the “noisy simulator” of the IBM Brisbane quantum computer and the “ideal simulator” of the noiseless IBM Brisbane quantum computer IBMsimulator1 ; IBMsimulator2 .
To perform the numerical experiments, we took the generator of the HQ-MolGAN-VVRQ model and separated it into two parts: VQC and MLP. In the experiments, we executed the VQC part on the simulator of the noisy and noiseless quantum hardware and fed the results to MLP executed on classical hardware. In the noisy simulation, we performed quantum operations on 8 noisy qubits with the best fidelities out of 127 available on the IBM Brisbane quantum computer. In the ideal simulation, we performed the same VQC with the same initial Gaussian-distributed vector on an ideal simulator.
The comparison between noisy and ideal simulations is shown in Fig. 5(e), where relative losses of noisy and ideal simulators with respect to the number of given shots are shown. The relative error is estimated as:
$\text{Err}(N)=\frac{|\sum_{n=1}^{256}[P_{\text{sim}\ i}(N)-P_{\text{ideal}\ i}(+\infty)]|}{\sum_{n=1}^{256}P_{\text{ideal}\ i}(+\infty)},$
where $P_{\text{ideal}}(+\infty)$ is the matrix of probabilities generated on the ideal simulator after a large number of iterations ($N\rightarrow+\infty$) and $P_{\text{sim}}(N)$ is the matrix of probabilities generated on the specific simulator (either ideal or noisy) after $N$ shots.
The relative error of the ideal simulator approaches zero as the number of shots grows, as shot noise is the only source of error. For the noisy simulator, the relative error hits a limit specific to the noise model of the quantum device. This systematic error of the device has an impact on the probability values given by the circuit, leading to slightly different initial states of the vector (Fig. 5(f)), which is given to the MLP layer in the course of generation.
[…]
@@ III Discussion
To enhance the classical MolGAN, we introduced two solutions: the incorporation of VQCs as the initial layer of the generator and the utilization of a cycle component to restore the original data from the graph representation of the generated molecule.
Our empirical results substantiate the merit of diversifying training datasets, not limiting to the QM9 dataset alone but also incorporating the PC9 dataset. Notably, the HQ-MolGAN model, with their Generator’s layers scaled down by half and trained for 4 times fewer iterations, has outperformed the classical MolGAN model decao2022molgan and its hybrid quantum analogues Li_Topaloglu_Ghosh_2021 ; Kao_2023 across key chemical metrics: QED, logP, SA, and uniqueness. The HQ-MolGAN model was also tested for potential execution on quantum computers. By using the noisy and ideal simulators of an IBM quantum computer, we observed that the HQ-MolGAN model is resilient to noise: the achieved scores of the noisy simulation are shown to be similar to the ideal noise-free simulation.
The introduction of the Cycle component to HQ-MolGAN, and especially its hybrid quantum variant, marks an advance in the training of hybrid quantum models. These models not only enhance the desired uniqueness score but also effectively mitigate the occurrence of the “high entropy state,” a notable challenge in molecular generation tasks. Consequently, these models hold substantial promise for applications in the domain of small drug compound design, both for commercial and scientific purposes within pharmacology.
@@
Looking ahead, we see a potential for hybrid quantum machine learning models to further advance the field of molecule generation using hybrid quantum models. We plan to delve deeper into refining the model architecture, particularly focusing on optimizing the balance between the quantum and classical components. This involves experimenting with different configurations and parameters to enhance the overall efficiency and accuracy of the models.
Another critical avenue we intend to pursue is the expansion of our training datasets. By incorporating a broader range of chemical compounds and molecular structures, we aim to increase the diversity and representativeness of our models. This expansion is expected to improve the models’ generalization capabilities and their applicability. Through these focused research efforts, we aspire to contribute significantly to the advancement of hybrid quantum computing in drug discovery, ultimately aiding in the development of more effective and innovative therapeutic solutions.
IV Supplementary Material
IV.1 Quantum Circuits Analysis
In this section, we analyze the quantum circuits employed in the HQ-Cycle-MolGAN framework, specifically the VVRQ and the QDI layers. We assess these circuits using several metrics:
  • ZX calculus circuit reducibility
  • Fisher information degeneracy
Figure 6: (a) VVRQ layer with original parameters. (b) VVRQ layer without redundant parameters. (c) QDI layer with original parameters. (d) QDI layer without redundant parameters.
IV.1.1 ZX calculus
ZX-calculus serves as a graphical language capable of depicting a quantum circuit through diagrams consisting of “spider”—nodes interconnected by edges. These ZX diagrams can be simplified duncan2022quantum and minimized using the language’s graphical rewriting rules wetering2020zx , which are grounded in the fundamentals of quantum operations. By simplifying these diagrams, we can derive a more efficient circuit configuration. Moreover, ZX-calculus offers a metric for evaluating circuit efficiency by comparing the number of parameters in the simplified diagram against the initial number of parameters. A reduction in redundant parameters indicates enhanced circuit performance. A circuit deemed unable to be optimized in this manner is classified as ZX-irreducible.
The key adjustments to the circuit shown in Fig. 6(a-b) consist of the rearrangement of certain weights after their reduction. During the optimization phase, $139$ out of $150$ parameters (approximately $93\%$) were preserved, illustrating the circuit’s significant degree of optimization. As illustrated in Fig. 6(c-d), for QDI, the ZX-calculus algorithm merely adjusted some of the weights following their reduction. Throughout the optimization process, $269$ out of $272$ parameters (about $99\%$) were maintained, indicating that the circuit is highly optimized and yields almost perfect outcomes.
By using the ZX-calculus algorithm, it was revealed that both VVRQ and QDI perform very well and have close to no redundant parameters. However, other metrics should be applied to obtain a more precise analysis.
IV.1.2 Fisher information
A supervised machine learning task can be described as the creation of a hypothesis model $h_{\theta}(\hat{x})$ based on a labeled dataset $(x,y)\in X\times Y$ to provide an approximation of the distribution, $f(x)$, of the data in nature. Using a subset of $S$ labeled data points from this distribution, we optimize our hypothesis model to provide high-accuracy modeling of $f(\hat{x})$. For this, we maximize the probability of acquiring the associated label $y$ from the model with parameters $\theta$ and data points $x$. The needed conditional probability can be written as $P(y|x,\theta)$. Taking into account the uniform distribution over $X$, the joint probability, $P(y,x|\theta)$ is used for better accuracy, and its distribution can be calculated for any value of $\theta$ for a data points $x_{i}$. Thus, we represent the joint probability as an N-dimensional manifold with $N$ as the number of trainable parameters $N=|\theta|$. The Fisher information matrix $F(\theta)$ abbas2020power ; amari1998gradient is a metric over this manifold.
$F(\theta)=\mathbb{E}_{\left\{x_{i},y_{i}\right\}}[\nabla_{\theta}\log(P)\nabla_{\theta}\log(P)^{T}]$ (1)
The next step is to diagonalize this metric to get a locally Euclidean tangential basis, where the diagonal values are the square gradient of our joint probability in this basis. These are the eigenvalues of the Fisher matrix. This is very important to detect and prevent the barren plateau problem, which involves vanishing gradients with a high number of qubits in quantum neural networks.
As was shown in Ref. mcclean2018barren , their expectation values become zero, and their variance decreases exponentially with a growing number of qubits. This can be seen if the gradients mostly degenerate near zero, which means that many parameters don’t participate in training at all. Therefore, calculating the eigenvalue spectrum of Fisher matrices for many realizations of $\theta$ helps investigate the trainability and robustness of the QNN against barren plateaus. A more highly trainable neural network would have less eigenvalue degeneracy.
The Fisher information matrix can be calculated for the specific hyperparameters of our circuit. Using a method from Ref. abbas2020power , we create a Gaussian dataset $x\sim\mathcal{N}(\mu=0,\sigma^{2}=1)$. Then the joint probability can be found by overlapping the computed state and the state of our quantum layer.
$P(y,x|\theta)=\braket{y}{\psi(\theta,x)},$ (2)
where $y$ is the output state. By averaging over all $x$ and $y$ we can calculate the Fisher information for any given $\theta$.
As a result, we can see in Fig. 7(a) that both circuits have at least half of their parameters significantly impacting the result. The QDI shows especially good results with four highly impactful parameters and two moderately impactful parameters. Fig. 7(b) shows the average Fisher matrices with no redundant elements on the diagonal. This shows that all the parameters are used in training, which leads to high trainability.
Ref. larocca2021overparam stated that some QNNs may show lowered parameter efficiency due to over-parametrization. This was calculated by finding that, at some point, parameter addition leaves the rank of the Fisher information matrix (FIM rank) unchanged. This happens when the circuit starts to become saturated. After that, there is no increase in expressivity, and there can be a risk of over-parametrization. As one can see in Fig. 7(d), the addition of new layers does not show any over-parametrization on this scale. Hence, the increase in rank with additional layers (and other metrics) can determine the necessity of structure change. As the previous analysis showed, VVRQ performs quite well, which may indicate that a more complex circuit is not needed. At least two layers are required since the addition of the second one changed the rank from $16$ to $38$ while the expected scenario would be an increase of the same amount as the first one (to $32$). It shows the underdevelopment of the first layer that is fixed by the addition of a new one. For QDI, the circuit is already optimized enough and does not show any underdevelopment for the first layer. Thus, the need for the addition of new layers can be determined by how much the rank increases, complemented by other methods (such as the Fisher eigenspectrum analysis).
Figure 7: (a) The normalized histogram of the VVRQ (left) and QDI (right) Fisher eigenspectrum. For VVRQ, all three layers have the first four parameters contributing the most impact (with the third layer adding extra frequency to the first parameter). This circuit is moderately expressive. For QDI, the majority of parameters achieved very high frequencies, which indicates excellent trainability. (b-c) Average Fisher matrices for VVRQ (left) and QDI (right). For VVRQ, the diagonal elements show that the circuit distributes the gradients to all trainable parameters with no evident single-parameter dominance (only a slight gradient shift towards the latter parameters for the deeper layers). For QDI, the diagonal elements show that the circuit distributes the gradients to all trainable parameters with no single-parameter dominance and almost no non-diagonal element, indicating high trainability. (d) FIM rank for VVRQ (left, $3/5$ layers used) and QDI (right, $1/5$ layers used) illustrates the circuit’s limit of over-parameterization. For VVRQ, the circuit isn’t over-parameterized, and the addition of the second layer more than doubled the rank (from $16$ to $38$). The third layer adds the same amount, so it’s not necessary to increase further. For QDI – taking into account excellent eigenspectrum performance and the usual increase from the first to the second layer (from $72$ to $144$, doubled, unlike the VVRQ), it’s unnecessary to add new layers.
IV.2 Hybrid Quantum Generator oscillatory behavior
HQ-MolGAN models have shown oscillatory behavior in terms of chemical metrics during training (Fig. 3). The reason for this phenomenon may lie in the breakdown of the interplay between the HQ-Generator and the classical Discriminator. To investigate this possibility, we provide a loss chart of the Discriminator component during the training of the classical MolGAN and the HQ-MolGAN-VVRQ (PC9).
Fig. 8 shows that the Discriminator competing with the HQ generator does not show any significant change compared to the classical MolGAN generator. Also, neither chart shows any significant increase after $15$ thousand iterations, which means that MolGAN’s generative expressiveness converges to a narrow beam of values due to the Generator’s properties and not the Discriminator’s.
Figure 8: Loss curve of the Discriminator for the classical MolGAN and the HQ-MolGAN-VVRQ.